Hi
i need help with this problem: (((6e^(2x)=18x^2))). i divided each side by x and took the ln of each side thus getting 2x=ln3x^2. now i don't know what to do. please help thanks

Answer by stanbon(75887)

6e^(2x)=18x^2)))
Divide both sides by 6 to get:
e^(2x)=3x^2
Graph the left side as y=e^(2x) and the right side as y=3x^2
and find the point(s) of intersection. You will have to have
a graphing calculator to get an accurate result.
I get x= -0.3906464...
Cheers,
Stan H.